# DyCE: Dynamic Configurable Exiting

### v1

@@ 7                   $\hat{\boldsymbol{y}}\leftarrow\tilde{\boldsymbol{y}}$
IV Design and Training
IV-A Exit Design
IV-A1 Network
The design of an exit must consider the trade-off between complexity and performance. Although more complex exit-ing functions have higher stand-alone performance, they would add more computational overhead and can affect the system’s overall performance. One of the possible forms of an exit is the multilayer perceptron (MLP)[20]. However, any functions that can generate a compatible output are eligible exiting functions. Meanwhile, since every exit is independent of others, there is no requirement for all exits to have the same architecture. They can be designed and trained separately. It is also possible to design multiple candidate exits for the same position. The searching algorithm (discussed in Sec. V) can find the most suitable one for different targets. In this paper, we use small MLPs as exits to generate results in the same (not just compatible as mentioned previously) format as $f_{N,1}$. We also employ an average pooling layer before the first MLP layer to reduce the feature map’s height and width to 1.
IV-A2 Feature Aggregation
@@ IV-B Training
IV-C Ensemble of Exits
The standalone performance of early attached exits might not be high because the preceding backbone segments (layers) are not sufficient in number and were not originally designed and trained to support this exit. However, by using a group of exits in a cooperative fashion, we significantly reduce incorrect predictions. Fig. 3 shows the accuracy of predictions at an early exit implemented on ResNet-34. As illustrated, fewer samples exit at higher thresholds but with better accuracies. This observation indicates that, though stand-alone accuracy at early exits may not be high enough, their high-confidence predictions (at higher thresholds) are very reliable. In a group of early-exits with high thresholds, samples must be exited by either a highly confident exit or by the final exit. This way, high overall accuracy can be maintained while reducing computational complexity for those samples which exit early. For varying the model performance in real-time, Sec. V discusses an algorithm to find the most suitable confidence range for every exit in such a way that the overall complexity is reduced for a specific performance target.
Figure 3: Accuracy of predictions at the $1^{st}$ exit (located at $1^{st}$ skip connection’s output) of ResNet-50 for ImageNet. As the threshold increases, fewer samples exit at this position while the accuracy is rising.
V Configuration Searching
@@ V-A Performance metrics
V-A1 Relative Overall Accuracy
@@
V-A2 Relative Run-time Complexity
@@
V-B Generating configurations
V-B1 Optimization Target
@@ V-B2 Searching for configurations
Here, $(k_{c},t_{c})$ denotes the latent action at the $n^{th}$ exiting position. If both $A_{\boldsymbol{k},\boldsymbol{t},n}(k_{c},t_{c})$ and $C_{\boldsymbol{k},\boldsymbol{t},n}(k_{c},t_{c})$ are monotonically increasing, then $f_{\boldsymbol{k},\boldsymbol{t},\lambda,n}(k_{c},t_{c})$ will be convex. When the threshold of any early exit is increased, more samples progress to the next exit, which invariably requires additional computation. Therefore, $C_{\boldsymbol{k},\boldsymbol{t},n}(k_{c},t_{c})$ is monotonically increasing. With regard to accuracy, as noted in Sec.IV-C, an increased threshold initially enhances the accuracy of the current exit. Moreover, samples reaching later exits are classified more accurately provided all exiting networks have a modest scale relative to the backbone, thus making $A_{\boldsymbol{k},\boldsymbol{t},n}(k_{c},t_{c})$ also monotonically increasing. Therefore, as long as later exits consistently require more computation and yield higher accuracy, $f_{\boldsymbol{k},\boldsymbol{t},\lambda,n}(k_{c},t_{c})$ is convex and its convergence is guaranteed.
Once $f_{\boldsymbol{k},\boldsymbol{t},\lambda,n}(k_{c},t_{c})$ can be minimized, we traverse different layers and exit types to ascertain the optimal action. This action is then applied to the existing configuration. This process is repeated until the round limit is reached, or no additional action can enhance the current configuration. This search method is encapsulated in our proposed Iterative Search, as depicted in Algo.2. Although this algorithm employs a greedy approach, which does not guarantee convergence to the optimum, it effectively narrows the search space to an acceptable range.
@@ 10                         $f_{min}\leftarrow f_{c}$
13      if $\mathrm{Candidate}~{}\text{is not}~{}\mathrm{None}$ then
@@ 15             $t_{n}\leftarrow t_{c}$       $k_{n}\leftarrow k_{c}$
where $\mathop{min}_{t_{c}}f_{\boldsymbol{k},\boldsymbol{t},\lambda,n}(k_{c},t_{c})$ is a bounded minimization algorithm to find a $t_{c}\in(0,1)$ such that the function value is minimized to $f_{c}$.
@@ 7                   $f_{min}\leftarrow f_{c}$       $t_{n}\leftarrow t_{c}$       $k_{n}\leftarrow k_{c}$
V-C Configurations for different targets
@@
VI Simulation
VI-A Base Model and Dataset
@@
In 2016, He et al. presented ResNet[14], which is a residual learning framework to train very deep networks for computer vision tasks. After that, many studies proved that ResNet and its residual structure have a successful performance on various computer vision tasks. The core element of ResNet is the residual block which consists of two or three cascaded convolutional layers with a shortcut. Many of today’s state-of-the-art deep CNNs still follow this structure. This paper validates the proposed early-exit based algorithm on the original version of ResNet[14] and ConvNextv2[22], but the results are applicable to other deep neural networks as well. ResNet was proposed for image classification based on ImageNet-1k[7] which contains 1.28m training images and 50k validation images with a size of 244x244. This paper simulates three variants of ResNet at different depths (ResNet-34, ResNet-50, and ResNet-152) on ImageNet to demonstrate the effectiveness of the proposed method.
VI-B Model Implementation
@@
The backbone part of our model implementation for the ImageNet dataset follows the repository by PyTorch[23]. We attach exits to the output of each residual block in ResNet. ResNet-34 and ResNet-50 have 16 positions for exits, including the last exit (the head of the original models). There are 50 exit points for ResNet-152 since it has more layers. The main purpose of this work is not to compete on performance; we employ five different types of plain MLP, from small to large, for every position except the last one. These five types are MLPs of 1-layer, 3-layer with 500 neurons in each hidden layer, 3-layer with 1000 neurons, 5-layer with 500 neurons, and 5-layer with 1000 neurons. The search algorithm will select either none or one of the options for each position in a configuration.
VI-B1 Training details
@@
The training of the exits is based on an existing pre-trained model. The backbone network is freezed and only the exits will be trained. The forward propagation will go through the whole model, while the backward propagation works only on each exit. The data augmentation follows the same strategy as ImageNet pre-training, and the optimizer is Adam with a learning rate $lr=3\times 10^{-4}$. The loss function is soft cross entropy as mentioned in Section IV-B.
VI-C Performance Evaluation
@@
In this study, we present an evaluation of performance to substantiate the effectiveness of our proposed method. Figure 4 provides a comparative visualization of the performance and computational requirements between our proposed approach and other methods. Our approach is validated with base models such as ResNet[14] and ConvNextv2[22]. Each point on the depicted curve signifies an available configuration, providing the option for the user to select in real-time.
TABLE II: Required Multiply-Accumulate operations (MACs) when a certain drop in accuracy (ACC) is allowed
[…]
  • 1
    Relative to the accuracy of base models
@@
Figure 4: MACs and Accuracy of the proposed method, base models and related work. Original ResNet and ConvNeXtv2 variants are represented as blue stars and orange triangles, the compressed version is denoted by the curve attached to each of them. Three related work is also plotted for reference. These models can not be configured in runtime or adaptive to new base models. However, DyCE can switch to different configurations (denoted by red dots) in runtime and adapt different models such as ResNet and ConvNextv2.
As a compression method, DyCE possesses the capability to conserve computational resources while incurring minimal performance degradation. Table II depicts the experimental results regarding the minimum Multiply-Accumulate operations (MACs) needed to retain specific accuracy levels. Table II depicts the experimental results regarding the minimum Multiply-Accumulate operations (MACs) needed to retain specific accuracy levels. Our proposed system requires only 76.5% to 79.4% of MACs to uphold a comparable accuracy (a diminution of $\leq 0.5\%$) for four variants of ResNet on the ImageNet dataset. As for the ConvNeXtv2 model, both the tiny and nano versions can be compressed by 18.1% and 25.9%, respectively, without significant accuracy degradation. However, the three smaller versions present a more challenging compression scenario, requiring a compromise of over 2% in accuracy for a similar reduction in MACs. For all these models, if a larger decrease in accuracy is permissible, DyCE has the potential to further curtail computational demands by approximately 15% to 20%, albeit with a 5% reduction in accuracy.
@@
Further information is visualized in Figure 4, where we have plotted the DyCE compression curve for each base model. Each dot on the curve denotes a run-time selectable configuration that enables a different trade-off point between complexity and performance. We employed a step size of 0.01 for $\lambda\in[0,1]$ to generate these curves. The density can be augmented to facilitate almost continuous performance tuning.
@@
DyCE exhibits advantages compared with the native architecture scaling, particularly in regions close to the base model. The DyCE curve for ResNet-152 initially surpasses the linear trajectory between ResNet-152 and ResNet-101 but subsequently falls slightly below the point representing ResNet-101. Nevertheless, ResNet-50 augmented with DyCE, consistently outperforms ResNet-34. These findings underscore the potential of DyCE as a beneficial supplement to conventional static model scaling.
@@
Comparative results from other dynamic scaling methods, such as MSDNet, MSNet, and SkipNet-101, are also depicted in Figure 4. These models are task-specific and not runtime configurable, their performance degrade more slowly compared to our plain MLP exits. Given the independence of each exit within DyCE, this challenge can potentially be mitigated by introducing dedicatedly designed exiting networks. Despite the simplicity of the MLPs, DyCE can surpass the efficiency of other methods when suitable base models are chosen. As the proposed framework is decoupled from the base model, it facilitates seamless integration with state-of-the-art models, such as ConvNextv2. The evaluation results of DyCE in conjunction with ConvNextv2 denote a substantial improvement over competing methods. While other techniques generally require model-specific knowledge for designing exits, they may not be easily applicable to the latest models. Considering the rapid evolution of deep learning, the flexibility offered by DyCE is a valuable feature.
VI-D Comparison between Searching Algorithms
@@
We have proposed two search algorithms for generating configurations: the iterative method, which is expected to yield better convergence, and the single-pass method, which offers significantly faster completion. The time complexity of the single-pass algorithm is directly proportional to the number of candidate exits, while the iterative method requires repetitive inspection of every exit, leading to substantially higher complexity. In our experiments, both algorithms have been implemented with GPU acceleration. The single-pass method requires 1.5 minutes to find 100 different configurations for ResNet-50, whereas the iterative method requires 28.9 minutes for completion. The disparity in time taken becomes more pronounced with an increase in the number of exits. For instance, ResNet-152, which has 50 exit positions, requires 4.9 minutes and 215.5 minutes to finish using the two respective methods.
@@
We also compared our proposed algorithms with a baseline method, which selects only one type of exit and applies identical thresholds to all exits. This is the most simplistic approach to make existing early-exit based systems configurable in real time. Figure 5 demonstrates that even when the best exit type is chosen (a 5-layer MLP with 1000 neurons in hidden layers), both of our proposed algorithms exhibit considerable advantages. These results indicate the effectiveness of the search algorithm we propose.
Figure 5: Comparison of three configuration generating methods on ResNet-50.
VI-E Ensemble vs Individual Exits
Figure 6 illustrates the efficacy of utilizing multiple exits collaboratively (represented by the purple line), demonstrating superior performance in comparison to individual exits (represented by the dots in dashed lines). To obtain standalone performance, we emulate a configuration with only one enabled exit and set its threshold to zero, ensuring inference consistently exits at that position. This process is repeated for all potential exits, culminating in the series of dots depicted. As a result, the most optimal type of exit network is “MLP-5x1000”, but the ensemble of exits significantly outperforms any individual types, including the optimal one. As discussed in Section.IV-C, the early exit in more trustable when a threshold is applied. In an exiting configuration, most exits except the final one have thresholds, making most exits perform better than their standalone performance. The systematic average computation is the weighted sum of every exit’s complexity but the accuracy is higher than the sum of exit’s accuracy with the same weight. This observation demonstrates the compression efficiency of early exit groups.

### v3

@@ 7                   $\hat{\boldsymbol{y}}\leftarrow\tilde{\boldsymbol{y}}$
4 Design and Training
4.1 Exit Design
4.1.1 Network
The design of an exit must consider the trade-off between complexity and performance. Although more complex exit-ing functions have higher stand-alone performance, they would add more computational overhead and can affect the system’s overall performance. One of the possible forms of an exit is the MLP. However, any functions that can generate a compatible output are eligible exiting functions. Meanwhile, since every exit is independent of others, there is no requirement for all exits to have the same architecture. They can be designed and trained separately. It is also possible to design multiple candidate exits for the same position. The searching algorithm (discussed in Sec. 5) can find the most suitable one for different targets. In this paper, we use small MLPs as exits to generate results in the same (not just compatible as mentioned previously) format as $f_{N,1}$. We also employ an average pooling layer before the first MLP layer to reduce the feature map’s height and width to 1.
4.1.2 Feature Aggregation
@@ 4.2 Training
4.3 Ensemble of Exits
The standalone performance of early attached exits might not be high because the preceding backbone segments (layers) are not sufficient in number and were not originally designed and trained to support this exit. However, by using a group of exits in a cooperative fashion, we significantly reduce incorrect predictions. Fig. 4 shows the accuracy of predictions at an early exit implemented on ResNet34. As illustrated, fewer samples exit at higher thresholds but with better accuracies. This observation indicates that, though stand-alone accuracy at early exits may not be high enough, their high-confidence predictions (at higher thresholds) are very reliable. In a group of early-exits with high thresholds, samples must be exited by either a highly confident exit or by the final exit. This way, high overall accuracy can be maintained while reducing computational complexity for those samples which exit early. For varying the model performance in real-time, Sec. 5 discusses an algorithm to find the most suitable confidence range for every exit in such a way that the overall complexity is reduced for a specific performance target.
Figure 4: Accuracy of predictions at the $1^{st}$ exit (located at $1^{st}$ skip connection’s output) of ResNet50 for ImageNet. As the threshold increases, fewer samples exit at this position while the accuracy is rising.
5 Configuration Searching
@@ 5.1 Performance metrics
5.1.1 Relative Overall Accuracy
@@
5.1.2 Relative Run-time Complexity
@@
5.2 Generating configurations
5.2.1 Optimization Target
@@ 5.2.2 Searching for configurations
Here, $(k_{c},t_{c})$ denotes the latent action at the $n^{th}$ exiting position (At $n^{th}$ position, exit $k_{c}$ is called to make a prediction and produce the confidence. The computing is terminated if the confidence is greater than $t_{c}$.). Empirically, the relationship between $A_{\boldsymbol{k},\boldsymbol{t},n}(k_{c},t_{c})$ and $C_{\boldsymbol{k},\boldsymbol{t},n}(k_{c},t_{c})$ is like Fig. 2. When more computation is used, the accuracy is improved and it will be more difficult to further improve the accuracy as the computation increases. i.e. We have:
$\frac{\partial A(k_{c},t_{c})}{\partial C(k_{c},t_{c})}>0~{}~{}$ $and~{}~{}\frac{\partial^{2}A(k_{c},t_{c})}{\partial C^{2}(k_{c},t_{c})}<0$ (14)
$\lim_{t_{c}\to 0}\frac{\partial A(k_{c},t_{c})}{\partial C(k_{c},t_{c})}=\infty~{}~{}$ $and~{}~{}\lim_{t_{c}\to 1}\frac{\partial A(k_{c},t_{c})}{\partial C(k_{c},t_{c})}=0$ (15)
That means, $\frac{\partial A_{\boldsymbol{k},\boldsymbol{t},n}(k_{c},t_{c})}{\partial C_{\boldsymbol{k},\boldsymbol{t},n}(k_{c},t_{c})}$ is positive and monotonically decreasing. Therefore, the equation $f^{\prime}_{\boldsymbol{k},\boldsymbol{t},\lambda,n}(k_{c},t_{c})=0$ has and only has one solution. i.e. $f_{\boldsymbol{k},\boldsymbol{t},\lambda,n}(k_{c},t_{c})$ has and only has one minimum. Hence, the object fuction is convex and the searching algorithm can converge to the global minimal.
Once $f_{\boldsymbol{k},\boldsymbol{t},\lambda,n}(k_{c},t_{c})$ can be minimized, we traverse different layers and exit types to ascertain the optimal action. This action is then applied to the existing configuration. This process is repeated until the round limit is reached, or no additional action can enhance the current configuration. This search method is encapsulated in our proposed Iterative Search, as depicted in Algorithm 2. Although this algorithm employs a greedy approach, which does not guarantee convergence to the optimum, it effectively narrows the search space to an acceptable range.
@@ 10                         $f_{min}\leftarrow f_{c}$
13      if Candidateis $\mathrm{Candidate}~{}\text{is not}~{}\mathrm{None}$ then
@@ 15             $t_{n}\leftarrow t_{c}$       $k_{n}\leftarrow k_{c}$
In Algorithm 2 $\mathop{min}_{t_{c}}f_{\boldsymbol{k},\boldsymbol{t},\lambda,n}(k_{c},t_{c})$ is a bounded minimization algorithm, e.g., golden section search to find a $t_{c}\in(0,1)$ such that the function value is minimized to $f_{c}$.
@@ 7                   $f_{min}\leftarrow f_{c}$       $t_{n}\leftarrow t_{c}$       $k_{n}\leftarrow k_{c}$
5.3 Configurations for different targets
@@
6 Simulation
6.1 Base Model and Dataset
@@
In 2016, He et al. presented ResNet[19], which is a residual learning framework to train very deep networks for computer vision tasks. After that, many studies proved that ResNet and its residual structure have a successful performance on various computer vision tasks. The core element of ResNet is the residual block which consists of two or three cascaded convolutional layers with a shortcut. Many of today’s state-of-the-art deep CNNs still follow this structure. This paper validates the proposed early-exit algorithm on ResNet [19], ConvNeXtv2 [28], and DaViT [29]. The observation is applicable to other deep neural networks as well. ImageNet-1k [16], a classic dataset for computer vision and artificial intelligence, contains 1.28 million training images and 50,000 validation images of size 224x224. It has been widely used as a benchmark for image classification, and results on the ImageNet-1k validation set are considered a reference for model generalizability in practice. This paper simulates variants of the three backbone architectures at different depths on ImageNet-1k to demonstrate the effectiveness of the proposed method.
6.2 Model Implementation
@@
The backbone part of our model implementation for the ImageNet dataset follows the repository by PyTorch[30]. We attach exits to the output of each residual block in the base model. For example, ResNet34 and ResNet50 have 16 positions for exits, including the last exit (the head of the original models). There are 50 exit points for ResNet152 since it has more layers. The main purpose of this work is not to compete on performance; we employ five different types of plain MLP, from small to large, for every position except the last one. These five types are MLPs of 1-layer, 3-layer with 500 neurons in each hidden layer, 3-layer with 1000 neurons, 5-layer with 500 neurons, and 5-layer with 1000 neurons. The search algorithm will select either none or one of the options for each position in a configuration.
6.2.1 Training details
@@
The training of the exits is based on an existing pre-trained model. The backbone network is freezed and only the exits will be trained. The forward propagation will go through the whole model, while the backward propagation works only on each exit. The data augmentation follows the same strategy as ImageNet pre-training, and the optimizer is Adam with a learning rate $lr=3\times 10^{-4}$. The loss function is soft cross entropy as mentioned in Sec. 4.2.
6.3 Performance Evaluation
@@
In this study, we present an evaluation of performance to substantiate the effectiveness of our proposed method. Fig. 5 provides a comparative visualization of the performance and computational requirements between our proposed approach and other methods. Our approach is validated with base models such as ResNet[19] and ConvNeXtv2[28]. Each point on the depicted curve signifies an available configuration, providing the option for the user to select in real-time.
Table 2: Comparison among DyCE enabled models and other models with configurable efficiency. The accuracy on the ImageNet validation set is reported. The DyCE enabled models are marked in bold.
[b] Model GMACs Acc DaViTbase[29] 15.98 84.49% ResNet152[19] 11.51 78.31% DyCE DaViTbase[29] 10.87 83.51% DaViTsmall[29] 8.77 84.01% ResNet101[19] 7.80 77.36% SkipNet[9] 6.65 77.61% DyCE DaViTsmall[29] 6.50 83.03% SkipNet[9] 5.64 76.61% DaViTtiny[29] 4.52 82.71% ConvNeXtV2tiny[28] 4.46 82.95% ResNet50[19] 4.09 76.14% CF-ViT[15] 3.98 81.85% SkipNet[9] 3.97 73.69% ResNet34[19] 3.66 73.30% DyCE ConvNeXtV2tiny[28] 3.27 82.03% CF-ViT[15] 3.13 80.87% DVT[14] 2.76 79.96% ConvNeXtV2nano[28] 2.45 81.87% SkipNet[9] 2.34 67.61% MSDNet[12] 2.08 74.59% DyCE ConvNeXtV2nano[28] 1.97 80.91% CF-ViT[15] 1.89 77.86% MSNet[13] 1.89 76.15% DVT[14] 1.73 78.98% MSDNet[12] 1.63 73.61% ConvNeXtV2pico[28] 1.37 80.32% DVT[14] 1.31 76.15% CF-ViT[15] 1.27 71.85% MSDNet[12] 1.23 70.62% MSNet[13] 1.20 75.16% DyCE ConvNeXtV2pico[28] 1.19 79.38% MSNet[13] 0.89 72.16% DVT[14] 0.88 70.06% ConvNeXtV2femto[28] 0.78 78.49% MSDNet[12] 0.78 64.76% DyCE ConvNeXtV2femto[28] 0.70 77.53% MSNet[13] 0.64 66.18% ConvNeXtV2atto[28] 0.55 76.64% DyCE ConvNeXtV2atto[28] 0.49 75.65%
@@
Figure 5: MACs and Accuracy of the proposed method, base models and related work. Original ResNet, ConvNeXtv2 and DaViT variants are represented as triangles, the compressed version is denoted by the curve attached to each of them. The performance curve of related work is also plotted for reference. These models can not be configured in runtime or adaptive to new base models. However, DyCE can switch to different configurations (denoted by red dots) in runtime and adapt models with different architectures.
DyCE, as a compression method, effectively conserves computational resources while maintaining minimal performance degradation. Tab. 2 compares DyCE-enabled models to other approaches. Compared to the original versions, DyCE reduces Multiply-Accumulate operations (MACs) by approximately 10% to 30% with only a small drop in accuracy. Furthermore, DyCE-enabled models outperform related methods by achieving higher accuracy within a comparable MAC range.
@@
Further information is visualized in Fig. 5, where we have plotted the DyCE compression curve for each base model. Each dot on the curve denotes a run-time selectable configuration that enables a different trade-off point between complexity and performance. We employed a step size of 0.01 for $\lambda\in[0,1]$ to generate these curves. The density can be augmented to facilitate almost continuous performance tuning.
@@
DyCE exhibits advantages compared with the native architecture scaling, particularly in regions close to the base model. The DyCE curve for ResNet152 initially surpasses the linear trajectory between ResNet152 and ResNet101 but subsequently falls slightly below the point representing ResNet101. Nevertheless, ResNet50 augmented with DyCE, consistently outperforms ResNet34. These findings underscore the potential of DyCE as a beneficial supplement to conventional static model scaling.
@@
Comparative results from other dynamic scaling methods, such as V-MoE, Mobile V-MoE, DVT, CF-ViT, MSDNet, MSNet, and SkipNet-101, are also depicted in Fig. 5. These models are task-specific and not runtime configurable, their performance degrade more slowly compared to our plain MLP exits. Given the independence of each exit within DyCE, this challenge can potentially be mitigated by introducing delicately designed exiting networks. Despite the simplicity of the MLPs, DyCE can surpass the efficiency of other methods when suitable base models are chosen. As the proposed framework is decoupled from the base model, it facilitates seamless integration with state-of-the-art models, such as ConvNeXtv2 and DaViT. The evaluation results of DyCE in conjunction with ConvNeXtv2 and DaViT denote a substantial improvement over competing methods. While other techniques generally require model-specific knowledge for designing exits, they may not be easily applicable to the latest models. Considering the rapid evolution of deep learning, the flexibility offered by DyCE is a valuable feature.
6.4 Comparison between Searching Algorithms
@@
We have proposed two search algorithms for generating configurations: the iterative method, which is expected to yield better convergence, and the single-pass method, which offers significantly faster completion. The time complexity of the single-pass algorithm is directly proportional to the number of candidate exits, while the iterative method requires repetitive inspection of every exit, leading to substantially higher complexity. In our experiments, both algorithms have been implemented with GPU acceleration. The single-pass method requires 1.5 minutes to find 100 different configurations for ResNet50, whereas the iterative method requires 28.9 minutes for completion. The disparity in time taken becomes more pronounced with an increase in the number of exits. For instance, ResNet152, which has 50 exit positions, requires 4.9 minutes and 215.5 minutes to finish using the two respective methods.
@@
We also compared our proposed algorithms with a baseline method, which selects only one type of exit and applies identical thresholds to all exits. This is the most simplistic approach to make existing early-exit based systems configurable in real time. Fig. 6 demonstrates that even when the best exit type is chosen (a 5-layer MLP with 1000 neurons in hidden layers), both of our proposed algorithms exhibit considerable advantages. These results indicate the effectiveness of the search algorithm we propose.
Figure 6: Comparison of three configuration generating methods on ResNet50.
6.5 Ensemble vs Individual Exits
Fig. 7 illustrates the efficacy of utilizing multiple exits collaboratively (represented by the purple line), demonstrating superior performance in comparison to individual exits (represented by the dots in dashed lines). To obtain standalone performance, we emulate a configuration with only one enabled exit and set its threshold to zero, ensuring inference consistently exits at that position. This process is repeated for all potential exits, culminating in the series of dots depicted. As a result, the most optimal type of exit network is “MLP-5x1000”, but the ensemble of exits significantly outperforms any individual types, including the optimal one. As discussed in Sec. 4.3, the early exit in more trustable when a threshold is applied. In an exiting configuration, most exits except the final one have thresholds, making most exits perform better than their standalone performance. The systematic average computation is the weighted sum of every exit’s complexity but the accuracy is higher than the sum of exit’s accuracy with the same weight. This observation demonstrates the compression efficiency of early exit groups.